# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ 5.3 The Potential of No-Gradient Unrolling
Existing studies focus on fixed tasks, i.e. pure prediction tasks when no differentiable solver is available, and correction tasks when employing a differentiable solver. Given the large and complex implementations of solvers used in current scientific projects, often with decade-long histories, approaches that do not require the differentiability of these numerical solvers are especially attractive. Our experiments quantify the expected benefits when switching from a prediction task to a correction task with a non-differentiable solver.
As illustrated in figure 15, this enables an average 4.5-fold accuracy improvement with the same network architecture and size. Small architectures can especially benefit from training with NOG utilizing a non-differentiable physical prior. The models with 0.5k parameters perform 9.36 and 3.83 times better when supported by a low-fidelity solver, where the second value originates from a deeper variant of the 0.5k parameterization. The factor remains relatively constant for the remaining sizes, which all deploy deep architectures. Medium-sized networks with 2k and 33k parameters see a slight decrease in these improvements with values of 2.80 and 3.23 respectively, while the largest networks’ accuracies improve by a factor of 3.82 and 3.96 once trained with NOG in a correction setup. In the KOLM case, we observe similar relative gains from introducing a numerical solver as physical prior. Here, the factor is 4.8 on average. Again, this relative factor is largest for small architectures, where we observe a value of 6.2 for the 32k network. Similarly to the KS case, the benefits of using a physical prior decrease for larger architectures. For our KOLM experiments the values are 4.19, 4.88, and 4.19 respectively. It has to be noted that these factors likely are specific to our setups and the numerical solvers chosen as a prior. However, the purpose of this evaluation is to illustrate the potential improvements made possible by correction unrolling. Crucially, only an interface between the solver and the network is sufficient for NOG correction training, the core of the solver can remain untouched. These experiments indicate that there could be large benefits if scientific codes were interfaced with machine learning (recommendation I, III). We believe that this training modality could be a new common ground for machine learning and the computational sciences. In the past, machine learning focused works often focused on purely predictive tasks. As our results indicate, an interface with existing numerical architectures is not only necessary to deploy hybrid models in a scientific computing environment but additionally unlocks further accuracy improvements when models are trained with NOG unrolling.
@@ 6 Conclusion
We have conducted in-depth empirical investigations of unrolling strategies for training neural PDE simulators. The inherent properties of unrolling were deduced from an extensive test suite spanning multiple physical systems, learning setups, network architectures, and network sizes. Our findings rendered five best practices for training autoregressive neural simulators via unrolling. Additionally, our test sets and differentiable solvers are meant to serve as a benchmark: The broad range of network sizes for popular architectures as well as the selection of common physical systems as test cases yield a flexible baseline for future experiments with correction and prediction tasks. We have identified non-differentiable unrolling as an attractive alternative to simplistic one-step training and implementations of differentiable solvers. Our results lay the foundation for assessing potential gains by deploying this method. Lastly, we measured convergence rates for neural networks and saw how these architectures converge poorly compared to numerical solvers. We concluded that deploying neural networks may thus be beneficial when a particular problem can benefit from their inherent properties, while grid scaling is still more efficient with classical approaches.
@@
In the future, problem-tailored network architectures could potentially improve the scaling of neural networks. However, at the current point of research, it is advisable to combine neural and numerical approaches in hybrid architectures to benefit from both. One of the applications that could benefit from this approach is training closure models for filtered or averaged PDE s. At the same time, there are limitations to our scope. Our physical systems live in the domain of non-linear chaos and are primarily connected to fluid mechanics. Consequently, there is no inherent guarantee that our results will straightforwardly transfer to different domains. Additionally, we mentioned the relevance of our results to the scientific computing community. Our recommendations can help prioritize implementation efforts when designing unrolled training setups. However, we have only studied a subset of the relevant hyperparameters. In the context of unrolling, variations such as network width and depth, advanced gradient-stopping techniques, and irregularly spaced curriculums could be impactful. These topics are promising avenues for future research to further our understanding of autoregressive neural networks for scientific applications.
@@ Acknowledgements
The code for the project is available at https://github.com/tum-pbs/unrolling.
@@
Appendix A Gradient Calculation
A.1 Correction
The network parameter optimization was defined in Equation 5. For one learning iteration, a loss was accumulated over an unrolled trajectory. We can write this total loss over the unrolled trajectory as a function of $g(\mathbf{u}^{i})=f_{\theta}(\mathcal{S}(\mathbf{u}^{i}))$ such that
$\mathcal{L}=\sum_{i=1}^{s}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{i+\tau s},g^{s}(\mathbf{u}^{i}))=\sum_{i=1}^{s}\mathcal{L}^{s},$ (14)
@@ A.2 Prediction
$\mathcal{L}=\sum_{i=1}^{s}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{i+\tau s},f_{\theta}^{s}(\mathbf{u}^{i}))=\sum_{i=1}^{s}\mathcal{L}^{s},$ (17)
@@
Appendix B AERO case
@@
$\mathcal{L}_{2}$ errors are computed and accumulated over 200 prediction steps (equivalent to 800 simulator steps) of simulation for two test cases with previously unseen Mach numbers: $M=0.725$ and $M=0.775$, corresponding to shock-free case and near-critical condition case.
Appendix C Additional Results
C.1 Prediction on the KOLM system
The networks used for the correction of the KOLM system were also trained for prediction. The results are visualized in figure 16. Unsurprisingly, the overall error values are larger than for the respective correction setups, where an additional numerical prior supports the network. Like the other prediction and correction setups, the errors decrease with increasing network size. Unrolled training modalities perform best at all network sizes. The quantitative values of the loss metric are also tabulated in 12.
@@ C.1 Prediction on the KOLM system
C.2 Interpolation and Extrapolation Tests
Interpreting physical hyperparameters is necessary for models to generalize to extrapolative test cases. We want to test whether extrapolation to new physical hyperparameters benefits from unrolling or long term gradients. In the main sections, all evaluations accumulated their results from interpolative and extrapolative test cases. In this section, we explicitly differentiate between interpolation and extrapolation tests. Figures 17, 18, 19, and 20 compare the model performance on interpolative and extrapolative test cases with respect to the physical parameters and depicts performance relative to the ONE baseline for various model sizes. Overall accuracy is worse on the harder extrapolative test cases. Similarly to the combined tests from the main section WIG performs best on average, both on interpolative or extrapolative data. However, the long-term gradients introduced by this method do not seem to explicitly favor inter- or extrapolation. Nonetheless, the positive aspects of WIG training are not constrained to interpolation, but successfully carry over to extrapolation cases.
@@ C.3 Dataset Sizes
C.4 Best performing models
@@ C.5 Additional Architecture
We tested the generalization of our findings toward attention-gated U-nets [44]. The implementation is a 1D version of the network used in the AERO cases introduced in Appendix B. Crucially, it features attention gates in its skip connections allowing the network to efficiently mix global and local structures in the output. The $\mathcal{L}_{2}$ errors of this architecture on the KS system are visualized in Figure 23. The behavior of ONE, NOG, and WIG matches well with other architectures (i.e. conv and graph nets) studied in the main section. We still observe the best performance with fully differentiable WIG unrolling, while NOG offers slimmer benefits over ONE. We can conclude that our observations regarding the positive effects of unrolling with and without gradients transfer to attention-based networks.
Appendix D Inference Visualizations
@@
Figure 24 displays a sample inference trajectory for convolutional correction models trained with the three unrolling variants. The inference case is selected from the extrapolation dataset, in this case for the small domain size of $\mathcal{X}=30.2$, which lies outside of the training range. The difference to the reference is further visualized in figure 25. Here, the distance to the target state as $u-\tilde{u}$ is used. This error was evaluated for the full range of model sizes as introduced in section 4 and listed in table 2. The same visualization procedure was also done for the other extrapolation test on large domains with $\mathcal{X}=70.4$. Figures 26 and 27 similarly display the state trajectories and errors for a sample network. A clear trend of longer correlation horizons is visible in the error figures. As already expanded on in the main paper, larger architectures stay in proximity to the target evolution for longer inference rollouts. At the same time, unrolling reduces errors for a fixed size since NOG and WIG have smaller amplitudes. On the large domain sizes in figures 26 and 27, the networks also show instabilities when trained with ONE or NOG.
For the KOLM system, we visualize sample rollout trajectories for the extrapolation case of Re=1000 in figure 28. The displayed horizon spans 250 timesteps. The final step of this inference rollout is further visualized in figure 29 for the whole range of training setups, including unrolling variants and network sizes. Finally, a visualization of the solution error is provided in figure 30. The visualizations show how those networks that were trained with ONE and NOG unrolling tend to misinterpret points of high vorticity and generally produce larger errors. As pointed out in the main sections, network size is, along with training modality, a decisive contributor to inference accuracy. This observation can also be made in the error visualizations in 30.
@@
Appendix E Evaluation Data

### v2

@@ 4.3 The Potential of No-Gradient Unrolling
Existing studies focus on fixed tasks, i.e. pure prediction tasks when no differentiable solver is available, and correction tasks when employing a differentiable solver. Given the large and complex implementations of solvers used in current scientific projects, often with decade-long histories, approaches that do not require the differentiability of these numerical solvers are especially attractive. Our experiments allow for quantifying the expected benefits when switching from a prediction task to a correction task with a non-differentiable solver.
As illustrated in Figure 19, this enables an average 13.3-fold accuracy improvement with the same network architecture and size. All architectures can benefit from training with NOG utilizing a non-differentiable numerical solver. The models with 0.5k parameters for the KS perform 23.8 and 17.5 times better when supported by a low-fidelity solver, where the second value originates from a deeper variant of the 0.5k parameterization. The factor remains relatively constant for the remaining sizes, which all deploy deep architectures. Medium-sized networks with 2k and 33k parameters see a slight decrease in these improvements with values of 14.3 and 25.7 respectively, while the largest networks’ accuracies improve by a factor of 30.0 and 18.9 once trained with NOG in a correction setup. In the KOLM case, we observe smaller relative gains from introducing a numerical solver. Here, the factor is 4.9 on average. This relative factor is largest for small architectures, where we observe a value of 6.4 for the 32k network. Similarly to the KS case, the benefits of using a numerical solver decrease for larger architectures. For our KOLM experiments the values are 4.19, 4.89, and 4.27 respectively. It has to be noted that these factors likely are specific to our setups and the used numerical solvers. However, the purpose of this evaluation is to illustrate the potential improvements made possible by correction unrolling. Crucially, only an interface between the solver and the network is sufficient for NOG correction training, the core of the solver can remain untouched. These experiments indicate that there could be large benefits if scientific codes were interfaced with machine learning (questions I, III). We believe that this training modality could be a new common ground for machine learning and the computational sciences. In the past, machine learning focused works often focused on purely predictive tasks. As our results indicate, an interface with existing numerical architectures is not only necessary to deploy correction models in a scientific computing environment but additionally unlocks further accuracy improvements when models are trained with NOG unrolling.
@@
4.4 Computational Performance
We measure the computational effort for both training and inference with our trained models. The ONE, NOG and WIG variants have different computational complexity when it comes to calculating a network update. ONE is the cheapest, as only one solver-network step needs to be computed. Both NOG and WIG are $m$-times more expensive in the forward pass. However, they differ in the backpropagation of gradients. The NOG setup does not differentiate solver operations, which makes the backpropagation pass cheaper than that of WIG. In principle, the gradient backpropagation of NOG setups is parallelizable across the unrolling horizon. While WIG requires a complete sequential backward pass through $m$ steps, NOG calculates independent gradient contributions for each unrolled step. We report numbers for the training of one GCN of size 1M on the KS system, as well as the CNN of size 0.1M on the KOLM system in Table 4(b). Here, it seems that the TensorFlow implementation of the KOLM setup is capable of taking advantage of this parallelization, reducing the training time of NOG significantly when compared to WIG.
At inference time, ONE, NOG, and WIG training variants all pose the same computational burden since their call signature is entirely identical. The computational time per inference step is listed in Table 4(a) for the correction networks of KOLM setup. Since the numerical procedure used in the KOLM setup is relatively expensive, the inference time is dominated by the numerical solver. Switching to deeper architectures only results in small additional computational costs.
Training time    $\text{KS}_{\text{graph}}^{\text{corr}}$    $\text{KS}_{\text{graph}}^{\text{pred}}$ $\text{KOLM}_{\text{conv}}^{\text{corr}}$
ONE 34 min. 28 min. 3.6 h
NOG 84 min. 64 min. 10 h
WIG 87 min. 67 min. 16 h
Inference time (s)
32k 0.0377
0.1M 0.0389
0.5M 0.0391
1M 0.0401
Table 4: (a): Computational time required for training a single graph model for KS of size 1M for correction and prediction, and a convolutional model for KOLM of size 0.1M trained for a correction task; (b): Inference time for a single timestep of the KOLM model in a correction setting, averaged across 100 steps
4.5 Discussion
Our work has taken a task- and architecture-agnostic stance. Combined with the large scale of evaluations across more than three thousand models, this allows for extracting general trends. In line with the research targets from Section 1, we arrive at the following set of recommendations:
(I) Non-differentiable unrolling as an alternative
The mitigating effects on data shift introduced by unrolling can boost model performance. An unrolled but non-differentiable chain can increase performance by 33% on average. In addition, non-differentiable neural correction based simulators do not require a new implementation of existing numerical solvers and thus represent an attractive alternative for fields with large traditional code bases. When no differentiable implementations are available, previous work typically resorts to pure prediction models. Instead, our results strongly encourage interfacing non-differentiable numerical implementations with machine learning, as doing so yields substantial accuracy improvements over prediction setups for identical network architectures and sizes.
(II) Differentiable unrolling delivers the best accuracy
When other hyperparameters are fixed, a differentiable unrolling strategy consistently outperforms other setups. At the cost of increased software engineering efforts, differentiable unrolling outperforms its one-step counterpart by 92% on average. Long-term gradients prove to be especially important for correction setups where neural networks correct numerical solvers.
(III) correction approaches yield improved scaling:
Our study across network sizes finds them predominant in terms of accuracy on inter- and extrapolative tests. However, we observe suboptimal convergence of the inference error when increasing the parameter count. These convergence rates compare poorly to the ones usually observed for numerical solvers. Most neural PDE simulators directly compete with numerical solvers. Thus, scaling to larger problems is critical for neural simulators. Unrolling can effectively train correction based simulators even when integrated with a non-differentiable solver. These combine the benefits of neural networks with the scaling properties of numerical solvers.
(IV) Low-dimensional problems match
While conducting these experiments, we noticed how training with multiple unrolled setups is non-trivial and sensitive to hyperparameters. Special care is thus required to successfully train with long unrolling. Reliable training setups utilize a curriculum where the unrolled number of steps slowly increases. The learning rate needs to be adjusted to keep the amplitude of the gradient feedback at a stable level.
@@ 5 Conclusion
We have conducted in-depth empirical investigations of unrolling strategies for training neural PDE simulators. The inherent properties of unrolling were deduced from an extensive test suite spanning multiple physical systems, learning setups, network architectures, and network sizes. Our findings rendered four best practices for training autoregressive neural simulators via unrolling. Most importantly, beneficial effects of unrolling at training time were present regardless of the differentiability of the neural simulator. We analyzed the effective unrolling horizons of the differentiable WIG and non-differentiable NOG setups, and brought these into context with the characteristic timescale of the physical system and the number of simulation steps. At the same time, we have shown how the unrolling horizons in the WIG cannot be derived from the chaotic timescale of the target dynamics alone. The target dynamics are inadequate to measure gradient behavior in the early stages of the network training, which motivated the curriculum approach. This behavior becomes even more complicated in the NOG case, where the temporal discretization affects the gradient stability. Deriving a general rule to choose ideal unrolling horizons in the WIG and NOG setups could be an interesting avenue for future research.
@@
Additionally, our test sets and differentiable solvers are meant to serve as a benchmark: The broad range of network sizes for popular architectures as well as the selection of common physical systems as test cases yield a flexible baseline for future experiments with correction and prediction tasks. We have identified non-differentiable unrolling as an attractive alternative to simplistic one-step training and implementations of differentiable solvers. Our results lay the foundation for assessing the gains that can be expected when deploying this method. Lastly, we measured convergence rates for neural networks and saw how these architectures converge poorly compared to numerical solvers. We concluded that deploying neural networks may thus be beneficial when a particular problem can benefit from their inherent properties, while grid scaling is still more efficient with classical approaches.
@@
In the future, problem-tailored network architectures could potentially improve the scaling of neural networks. However, at the current point of research, it is advisable to combine neural and numerical approaches in correction setups to benefit from both. One of the applications that could benefit from this approach is training closure models for filtered or averaged PDE s, although non-differentiable unrolling in this context is restricted to correction-style formulations. At the same time, there are limitations to our scope. Our physical systems live in the domain of non-linear chaos and are primarily connected to fluid mechanics. Consequently, there is no inherent guarantee that our results will straightforwardly transfer to different domains. Additionally, we mentioned the relevance of our results to the scientific computing community. Our recommendations can help prioritize implementation efforts when designing unrolled training setups. However, we have only studied a subset of the relevant hyperparameters. In the context of unrolling, variations such as network width and depth, advanced gradient-stopping techniques, and irregularly spaced curriculums could be impactful. These topics are promising avenues for future research to further our understanding of autoregressive neural networks for scientific applications.
@@ Acknowledgements
The code for the project will be made available at https://github.com/tum-pbs/unrolling.
@@
Appendix A Unrolling on Attractors
Here we provide a proof for equation (10) for a one-dimensional system. It is based on the definitions and remarks shown in Chapter 4 of [12]. The discrete equivalent of an $\omega$-limit set defined therein is
$\omega(u)=\{y\in\mathbb{R}|\forall T\in\mathbb{N}\hskip 5.0pt\exists t>T:g^{t}(u)=y\}.$ (35)
The $\omega$-limit set of an initial $u$ consists of the accumulation states that the trajectory converges to for large time horizons. In the following, we will use the $\omega$-limit set to show how unrolling converges to the learned dynamics attractor. Furthermore, it is shown that $w(u)$ is an attractor of the dynamics $A_{g}$ if there is an arbitrarily small neighborhood $\mathcal{N}_{u}$ such that
$g^{t}(\mathcal{N}_{u})\subset\omega(u)\hskip 5.0pt\forall t\hskip 5.0pt\text{and}\hskip 5.0pt\bigcap_{t\geq 0}g^{t}(\mathcal{N}_{u})=\omega(u)=A_{g}.$ (36)
Lemma
Consider a learned simulator $g:\mathbb{R}\xrightarrow{}\mathbb{R}$ being autoregressively unrolled at training for $m$ steps. Suppose that an attractor $A_{g}$ exists for the dynamics of $g$, and that the system is unrolled from states $u$ in a dataset $U$. The unrolling then converges to this attractor for large $m$ and sufficiently many initial states in the basin $u\in\mathcal{B}(A_{g})$.
Proof: An unrolled trajectory consists of $\mathrm{traj}(u)=\{g^{s}(u)|s\in[0,m]\}$. As $m\xrightarrow{}\infty$, the trajectory includes the $\omega$-limit set such that $\omega(u)\subset\mathrm{traj}(u)$. Then, with sufficiently many initial states in the basin $u\in\mathcal{B}(A_{g})$ and using (36)
$\bigcap_{u}\mathrm{traj}(u)\supset\omega(u)=A_{g}$ (37)
Appendix B Gradient Calculation
B.1 Correction
The network parameter optimization was defined in Equation 11. For one learning iteration, a loss was accumulated over an unrolled trajectory. We can write this total loss over the unrolled trajectory as a function of $g(\mathbf{u}^{t})=f_{\theta}(\mathcal{S}(\mathbf{u}^{t}))$ such that
$\mathcal{L}=\sum_{t=1}^{s}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{t+\tau s},g^{s}(\mathbf{u}^{t}))=\sum_{t=1}^{s}\mathcal{L}^{s},$ (38)
@@ B.2 Prediction
$\mathcal{L}=\sum_{i=t}^{s}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{t+\tau s},f_{\theta}^{s}(\mathbf{u}^{t}))=\sum_{t=1}^{s}\mathcal{L}^{s},$ (41)
@@
Appendix C AERO case
@@
We evaluate the vortex shedding frequency to identify the characteristic timescale of the AERO simulations. The frequency $f_{s}c/U_{\infty}$ was individually evaluated for every hyper-parameterization in the training dataset and averaged to obtain a characteristic timescale $t_{c,\text{AERO}}=1/{f_{s}}=0.455$. This timescale is covered by roughly 15 prediction steps of the neural simulator.
@@
$\mathcal{L}_{\mathrm{rel}}$ errors are computed and accumulated over 13 prediction steps (equivalent to 52 simulator steps) of simulation for two test cases with previously unseen Mach numbers: $M=0.725$ and $M=0.775$, corresponding to shock-free case and near-critical condition case.
Appendix D Additional Results
D.1 Metric Comparison
The choice of metric can affect the interpretation of results. In the main evaluations of the paper, we showcase the results in terms of a relative error metric, which is normalized by ground truth data. Relative error values approaching 1 indicate that the compared trajectory de-correlates from the ground truth. This effect is especially pronounced in chaotic dynamics and motivates shorter evaluation horizons (e.g., 12 steps for KS). As a long-term stability measure, we used the divergence-time in figure 9. By putting a threshold on the $\mathcal{L}_{2}$ error, this metric evaluates the number of steps for which the solution remains stable. However, it does not contain any information about the accuracy of a solution with respect to the ground truth trajectory. The absolute $\mathcal{L}_{2}$ metric in figure 20 can be seen as an intermediate between relative and divergence-time metrics. It depicts both the accuracy of a solution compared to the ground truth and the growing energies in divergent trajectories. In the present evaluation of the KS system, we have thus used intermediate-length horizons (40 steps).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 20: Metric comparison between relative and absolute error metrics
D.2 Prediction on the KOLM system
The networks used for the correction of the KOLM system were also trained for prediction. The results are visualized in Figure 21. Unsurprisingly, the overall error values are larger than for the respective correction setups, where an additional numerical prior supports the network. Like the other prediction and correction setups, the errors decrease with increasing network size. Unrolled training modalities perform best at all network sizes. The quantitative values of the loss metric are also tabulated in 13.
@@ D.2 Prediction on the KOLM system
D.3 Interpolation and Extrapolation Tests
Interpreting physical hyperparameters is necessary for models to generalize to extrapolative test cases. We want to test whether extrapolation to new physical hyperparameters benefits from unrolling or long-term gradients. In the main sections, all evaluations accumulated their results from interpolative and extrapolative test cases. In this section, we explicitly differentiate between interpolation and extrapolation tests. Figures 22, 23, 24, and 25 compare the model performance on interpolative and extrapolative test cases with respect to the physical parameters and depicts performance relative to the ONE baseline for various model sizes. Overall accuracy is worse on the harder extrapolative test cases. Similarly to the combined tests from the main section WIG performs best on average, both on interpolative or extrapolative data. However, the long-term gradients introduced by this method do not seem to explicitly favor inter- or extrapolation. Nonetheless, the positive aspects of WIG training are not constrained to interpolation, but successfully carry over to extrapolation cases.
@@ D.4 Dataset Sizes
D.5 Best performing models
@@ D.6 Additional Architecture
We tested the generalization of our findings toward attention-gated U-nets [62]. The implementation is a 1D version of the network used in the AERO cases introduced in Appendix C. Crucially, it features attention gates in its skip connections allowing the network to efficiently mix global and local structures in the output. The $\mathcal{L}_{2}$ errors of this architecture on the KS system are visualized in Figure 28. The behavior of ONE, NOG, and WIG matches well with other architectures (i.e. conv and graph nets) studied in the main section. We still observe the best performance with fully differentiable WIG unrolling, while NOG offers slimmer benefits over ONE. We can conclude that our observations regarding the positive effects of unrolling with and without gradients transfer to attention-based networks.
Appendix E Inference Visualizations
@@
Figure 29 displays a sample inference trajectory for convolutional correction models trained with the three unrolling variants. The inference case is selected from the extrapolation dataset, in this case for the small domain size of $\mathcal{X}=30.2$, which lies outside of the training range. The difference to the reference is further visualized in Figure 30. Here, the distance to the target state as $u-\tilde{u}$ is used. This error was evaluated for the full range of model sizes as introduced in Section 3 and listed in Table 2. The same visualization procedure was also done for the other extrapolation test on large domains with $\mathcal{X}=70.4$. Figures 31 and 32 similarly display the state trajectories and errors for a sample network. A clear trend of longer correlation horizons is visible in the error figures. As already expanded on in the main paper, larger architectures stay in proximity to the target evolution for longer inference rollouts. At the same time, unrolling reduces errors for a fixed size since NOG and WIG have smaller amplitudes. On the large domain sizes in Figures 31 and 32, the networks also show instabilities when trained with ONE or NOG.
For the KOLM system, we visualize sample rollout trajectories for the extrapolation case of Re=1000 in Figure 33. The displayed horizon spans 250 timesteps. The final step of this inference rollout is further visualized in Figure 34 for the whole range of training setups, including unrolling variants and network sizes. Finally, a visualization of the solution error is provided in Figure 35. The visualizations show how those networks that were trained with ONE and NOG unrolling tend to misinterpret points of high vorticity and generally produce larger errors. As pointed out in the main sections, network size is, along with training modality, a decisive contributor to inference accuracy. This observation can also be made in the error visualizations in 35.
@@
Appendix F Evaluation Data